# Neutrino Oscillations in Core-Collapse Supernovae and Neutron Star Mergers

### v1

@@ Abstract
Accurate neutrino transport is crucial for reliably modeling explosive astrophysical events like core-collapse supernovae (CCSNe) and neutron star mergers (NSMs). However, in these extremely neutrino-dense systems, flavor oscillations exhibit challenging nonlinear effects rooted in neutrino–neutrino forward scattering. Evidence is quickly accumulating that these collective phenomena can substantially affect explosion dynamics, neutrino signals, nucleosynthesis, and kilonova light curves. We review the progress made so far on the difficult and conceptually deep question of how to correctly include this physics in simulations of CCSNe and NSMs. Our aim is to take a broad view of where the problem stands, and so provide a critical assessment of where it is headed.
@@ 1 INTRODUCTION
Neutrino transport theory becomes much less settled when we shift our attention to the more extreme stellar environments of core-collapse supernovae (CCSNe) and neutron star mergers (NSMs). These are two events of great importance in the era of multimessenger astronomy, but we do not yet know how to correctly incorporate oscillations into the relevant predictions. The growing evidence that oscillations meaningfully alter the modeling of these sites has moved this open question—what we call the oscillation problem—to the forefront of theoretical astrophysics.
@@
The focus of this review is the intense ongoing effort to solve the oscillation problem. Our objective is to provide a useful guide, not an exhaustive one. Classic reviews from 2010 [1] and 2016 [2] remain valuable resources on collective neutrino oscillations. A recent comprehensive review [3] touches on a number of topics largely omitted here, and several others particularly offer greater coverage of nucleosynthesis [4], flavor instabilities [5, 6, 7], many-body correlations [8], and the early universe [9].
@@ 2.2 Prevalence of flavor instabilities
Neutrino oscillations introduce mesoscale dynamics. Linear stability analysis [16, 17] allows us to begin addressing the obvious follow-up question: How important is it? Flavor instabilities that appear in post-processed simulation data are evidence that oscillations, if they had been included in the simulation, would have caused some amount of deviation from the results obtained without them. A complete picture of the prevalence of instabilities is still being developed, but already we can see that they are widespread.
A turning point in this research area was the realization that the conditions for fast flavor instability (FFI) [18, 19] are met over a range of times and locations relevant for CCSN and NSM observables [20, 21, 22, 23, 24, 25, 26, 27] (Fig. 2). FFI occurs when the angular distributions of $\nu_{e}$ and $\bar{\nu}_{e}$ are sufficiently distinct. The angular distributions of all species are nearly isotropic in the neutrino trapping region and become increasingly forward-peaked as a function of distance. There is a species-dependence overlaid on this trend that has two principal origins. Firstly, charged-current processes involving nucleons generate imbalances that reflect the neutron-to-proton ratio. For example, in neutron-rich environments, neutrino absorption on nucleons delays the decoupling of $\nu_{e}$ relative to $\bar{\nu}_{e}$. Secondly, neutral-current scattering acts more strongly on $\bar{\nu}_{e}$ than $\nu_{e}$ due to the former’s higher average energy. These factors tend to deplete $\nu_{e}$ in outgoing directions and enhance $\bar{\nu}_{e}$ in ingoing directions. The result is that FFI is pervasive in multidimensional simulations, with an associated length scale commonly on the order of centimeters.
Evidence indicates that collisional flavor instability (CFI) [28] is also prevalent [29, 30, 31, 32, 33, 34]. This type of instability results from sufficiently distinct collision rates of $\nu_{e}$ and $\bar{\nu}_{e}$. The prime locations for CFI are the neutrino decoupling regions. In trapping regions, flavor instabilities of all kinds tend to be suppressed by neutrino degeneracy. In free-streaming regions, collisional rates are so low that CFI is probably unimportant even where it does appear. But in between, CFI is common and typically grows on a length scale of $\mathcal{O}(10)$ km. This scale can be much smaller when CFI is resonant [30, 35].
@@ 2.3 Estimated effects of oscillations
Figure 3: Potential astrophysical effects of neutrino oscillations. Left: Angle-averaged shock radius in 2D CCSN simulations with flavor conversion setting in above different critical densities (color-coded, with labels in g cm-3), showing significant deviations from the simulation without oscillations (“noFC”). Adapted with permission from Ehring et al. (2023) [49]. Middle: Nucleosynthesis abundance distributions formed in the simulated ejecta of a post-merger accretion disk, for various flavor-mixing prescriptions. Enhanced lanthanide abundances are observed. Modified with permission from Just et al. (2022) [50]. Right: CCSN neutrino energy spectra in various possible flavor-mixing scenarios, with potentially observables consequences. Adapted with permission from Abbar and Volpe (2024) [51].
@@
Flavor instabilities open the door to much more substantial effects on the astrophysics. For instance, as we saw in Sec. 2.2, flavor conversion likely occurs behind the supernova shock. Estimates indicate that explosion dynamics could be significantly affected [66, 49, 67, 68]. The left panel of Fig. 3 exemplifies the findings obtained using approximate treatments of oscillations in CCSN simulations. Notably, the evolution of the shock is highly variable depending on precisely where the flavor conversion sets in. This observation suggests that rough approximations, though clearly valuable, are not an adequate answer to the oscillation problem. The details matter.
Flavor instabilities might augur significant changes to nucleosynthesis [69, 70, 26, 50, 71] (middle panel of Fig. 3). For example, flavor conversion in a CCSN may substantially affect neutrino-induced nucleosynthesis in outer stellar envelopes [62, 64]. A strong influence is also possible in the outflows from post-merger accretion disks, which are sensitive to neutrinos emanating from the accretion disk and the central neutron star if one exists. In this case, flavor conversion alters the electron fraction in the disk outflow and affects the yield of heavy neutron-rich elements [72, 26, 50, 27]. The modification of nuclide abundances could in turn appreciably affect the kilonova light curve produced by the decay of radioactive elements [26, 50]. Firming up the quantitative connection between flavor conversion and kilonovae is a high priority for this research area.
@@ 3.1 Quantum coherence and Wigner functions
However, solving for the complete quantum evolution is inconceivable from a practical standpoint. The spirit of quantum kinetics is to be selective in retaining and evolving information characterizing the field. Expectation values of the form $\langle a^{\dagger}(t)a(t)\rangle$ are generalized occupation numbers, with the generalization referring to the fact that $a^{\dagger}$ and $a$ may differ in their mass states, helicities, or momenta. In such cases, and in the chosen basis, $\langle a^{\dagger}(t)a(t)\rangle$ is not a density of particles per se but rather a density of quantum coherence. This review is concerned with mass coherence, or nonzero expectation values of the type $\langle a^{\dagger}_{j}(t)a_{i}(t)\rangle$ for $j\neq i$. Helicity coherence, $\langle a^{\dagger}_{h^{\prime}}(t)a_{h}(t)\rangle\neq 0$ for $h^{\prime}\neq h$, and opposite-momentum pairing, $\langle b(t,-\bm{q})a(t,\bm{q})\rangle\neq 0$, are generally suppressed by the high scale of neutrino kinetic energy [74, 75, 76, 77].
Some amount of momentum coherence must be tolerated to allow for spatial inhomogeneity. To this end, the Wigner function (often called the density matrix) is introduced [78, 79]:
@@ 3.2 Coherent and incoherent processes
The commutator $[H_{\bm{q}},\rho_{\bm{q}}]$ accounts for coherent processes [10, 11, 12, 13, 1]:
@@ 3.3 Many-body correlations
Several challenges will have to be met as the quantum many-body formulation of neutrino transport continues to develop. We emphasize two that are specific to this problem. The first originates from the fact that Eq. 6 is actually a requantization of the kinetic Hamiltonian (Eq. 4) and, as such, unphysically privileges forward scattering [90]. At the time of writing, only one study has carried out calculations using the complete many-body Hamiltonian [91]. It reported outcomes qualitatively different from those calculated using Eq. 6, observing flavor equilibration and momentum thermalization on the same time scale.
The second challenge is that calculations need to evolve appropriate initial states. In particular, the applicability of quantum kinetics may hinge on the existence of localized particles [92, 90]. All fully many-body calculations performed so far have evolved states that are initially product states in the basis of definite-$\bm{q}$ plane waves. It is unclear whether kinetic evolution should be expected for such systems. The quasi-many-body model of Ref. [93] highlights this point. Two-body correlations build up during each brief interaction but are then traced out, following the logic of molecular chaos. The evolution in this model exhibits distinct coherent and incoherent contributions, as expected of a kinetic system. An important goal for this research area is to definitively establish the relationship between kinetic evolution and the properties of the underlying quantum state or ensemble.
@@ 4.1 Symmetries and conservation laws
An advantage of polarization vectors is that their motion is visualizable. Consider the QKE for $\bm{P}_{\bm{q}}$, which follows from plugging the definition above into Eq. 3:
@@
with $\omega_{\bm{q}}=\delta m^{2}/2|\bm{q}|$ and $\delta m^{2}=m_{2}^{2}-m_{1}^{2}$. The signs of $\delta m^{2}$ and $\omega_{\bm{q}}$ reflect the neutrino mass hierarchy. All flavor-space vectors are defined through expansions similar to Eq. 8. We write $\bm{\vec{D}}_{1}$ with an arrow to distinguish it from $\bm{D}_{1}$ in Eq. 5. By our chosen $\bar{\rho}$ convention, $\bm{\bar{H}}_{\bm{q}}$ is identical to $\bm{H}_{\bm{q}}$ but with $\omega\bm{B}\rightarrow-\omega\bm{B}$. We will stick with the flavor basis from this point on, so that $\bm{z}$ always represents the flavor axis. Assuming the muon and tau populations are negligible, the charged-lepton vector is $\bm{L}=n_{e}\bm{z}$, with net electron rest number density $n_{e}$. The mass vector is $\bm{B}=(\sin 2\theta)\bm{x}-(\cos 2\theta)\bm{z}$, where $\theta$ is the vacuum mixing angle.
@@
A defining feature of the oscillation problem is that calculations are only tractable for relatively simple models. Simplification, however, is an art. Often it introduces symmetries and additional invariants that can artificially stabilize neutrino flavor states and otherwise fundamentally change the dynamics, as strikingly demonstrated by the slow [98, 99, 100, 101], fast [102, 103, 100, 101], and collisional [104] flavor pendula. The potential for insight is balanced by the risk of unwarranted extrapolation. With all models, it is wise to ask which features are general, which are informatively specific, and which are just irrelevant.
@@ 4.2 Adiabaticity
One of the puzzles to come out of the bulb model was that the adiabatic approximation sometimes worked and sometimes did not. The effort to determine the conditions under which oscillations are adiabatic motivated the formalization of linear stability analysis as it applies to neutrino flavor (Sec. 4.3). Ultimately, the systematic study of flavor instabilities did not merely clarify the evolution that had been observed numerically. It undermined the use of the bulb model as a way to predict flavor transformation in the first place. The reason was already alluded to in Sec. 4.1: the stringent assumptions of the bulb model restrict the evolution in unacceptable ways, as evidenced by the presence of symmetry-breaking flavor instabilities. The story for NSMs parallels the one for CCSNe. In simple merger models, adiabaticity explains the numerically observed evolution through matter–neutrino resonances [57, 58, 111, 112, 113, 114]. These models appear to be compromised by oversimplification as well [115, 116].
@@ 4.3 Instability
By itself, this procedure does not shed much light on what gives rise to flavor instabilities. Deeper understanding comes from considering the linearized evolution of a neutrino system in certain idealized limits. Here the common three-type classification of instabilities enters (Sec. 2.2). The first class is FFI [18, 19]. By definition, an instability is fast if it occurs when all vacuum oscillation frequencies $\omega$ and all collision rates $\Gamma$ are set to zero in the dispersion relation. Growth rates can be as high as $\mathcal{O}(\mu)$, where $\mu=\sqrt{2}G_{F}n_{\nu}$. The category of SFI includes all the collisionless instabilities that disappear at $\omega=0$ [36, 37]. The speediest among them have $\mathcal{O}(\sqrt{\mu\omega})$ growth rates. CFI rounds out the taxonomy [28]. As the name implies, they disappear at $\Gamma=0$. Their growth rates reach up to $\mathcal{O}(\sqrt{\mu\Gamma})$. It is important to stress that these rates are highly approximate and are not even the correct scaling in all cases. For example, FFI identified in a simulation usually has a rate quite a bit smaller than $\mu$, and CFI more typically has a rate roughly proportional to $\Gamma$.
Occasionally one encounters other instability types, such as bipolar, multi-azimuthal-angle, and multi-zenith-angle. These terms are somewhat more historical, though they can still be useful depending on the context. For the newcomer to collective oscillations, the important thing to know is that they do not refer to distinct phenomena not covered here. The bipolar instability, for instance, is a particular manifestation of SFI.
@@
The concept of self-induced resonance helps illuminate the physics of flavor instabilities [126, 102]. We will explain the basic idea using the equation of motion $\partial_{t}\bm{P}_{\bm{q}}=\bm{H}_{\bm{q}}\times\bm{P}_{\bm{q}}$. Each polarization vector $\bm{P}_{\bm{q}}$ attempts to precess around its Hamiltonian $\bm{H}_{\bm{q}}$, but $\bm{H}_{\bm{q}}$ itself is evolving. If the relative azimuthal angle between these vectors evolves rapidly, then the term $\bm{H}_{\bm{q},T}\times\bm{P}_{\bm{q},T}$, which drives the evolution of $P_{\bm{q},z}$, is expected to average to zero over time. However, if $\bm{H}_{\bm{q},T}$ and $\bm{P}_{\bm{q},T}$ maintain a fixed relative phase over a long enough duration—in other words, if there is a resonance between the frequencies—then a secular change in $P_{\bm{q},z}$ can accumulate. This idea extends to spatially inhomogeneous systems, where individual neutrinos can additionally be resonant with $\bm{K}\neq 0$ collective waves [120].
The resonance mechanism is somewhat different in CFI [28, 104]. Consider the idealized system governed by $\partial_{t}\bm{P}=\sqrt{2}G_{F}\bm{D}\times\bm{P}-\Gamma\bm{P}_{T}$ and $\partial_{t}\bm{\bar{P}}=\sqrt{2}G_{F}\bm{D}\times\bm{\bar{P}}-\bar{\Gamma}\bm{\bar{P}}_{T}$, which are coupled through $\bm{D}=\bm{P}-\bm{\bar{P}}$. Introducing $\Gamma^{\pm}=(\Gamma\pm\bar{\Gamma})/2$, we rewrite these equations in terms of the difference vector $\bm{D}$ and sum vector $\bm{S}=\bm{P}+\bm{\bar{P}}$:

### v2

@@ Abstract
Accurate neutrino transport is crucial for reliably modeling explosive astrophysical events like core-collapse supernovae (CCSNe) and neutron star mergers (NSMs). However, in these extremely neutrino-dense systems, flavor oscillations exhibit challenging nonlinear effects rooted in neutrino–neutrino forward scattering. Evidence is quickly accumulating that these collective phenomena can substantially affect explosion dynamics, neutrino and gravitational-wave signals, nucleosynthesis, and kilonova light curves. We review the progress made so far on the difficult and conceptually deep question of how to correctly include this physics in simulations of CCSNe and NSMs. Our aim is to take a broad view of where the problem stands, and so provide a critical assessment of where it is headed.
@@ 1 INTRODUCTION
Neutrino transport theory becomes much less settled when we shift our attention to the more extreme stellar environments of core-collapse supernovae (CCSNe) and neutron star mergers (NSMs). These are two types of events of great importance in the era of multimessenger astronomy, but we do not yet know how to correctly incorporate oscillations into the relevant predictions. The growing evidence that oscillations meaningfully alter the modeling of these sites has moved this open question—what we call the oscillation problem—to the forefront of theoretical astrophysics.
@@
The focus of this review is the intense ongoing effort to solve the oscillation problem. Another review in the same volume of this journal covers other aspects of supernova modeling [1]. Our objective is to provide a useful guide, not an exhaustive one. Classic reviews from 2010 [2] and 2016 [3] remain valuable resources on collective neutrino oscillations. A recent comprehensive review [4] touches on a number of topics largely omitted here, and several others particularly offer greater coverage of nucleosynthesis [5], flavor instabilities [6, 7, 8], many-body correlations [9], the early universe [10], and neutrino astronomy [11].
@@ 2.2 Prevalence of flavor instabilities
Neutrino oscillations introduce mesoscale dynamics. Linear stability analysis [18, 19] allows us to begin addressing the obvious follow-up question: How important are oscillations? Flavor instabilities that appear in post-processed simulation data are evidence that oscillations, if they had been included in the simulation, would have caused some amount of deviation from the results obtained without them. A complete picture of the prevalence of instabilities is still being developed, but already we can see that they are widespread.
A turning point in this research area was the realization that the conditions for fast flavor instability (FFI) [20, 21] are met over a range of times and locations relevant for CCSN and NSM observables [22, 23, 24, 25, 26, 27, 28, 29] (Fig. 2). FFI occurs when the angular distributions of $\nu_{e}$ and $\bar{\nu}_{e}$ are sufficiently distinct. The angular distributions of all species are nearly isotropic in the neutrino trapping region and become increasingly forward-peaked as a function of distance from the principal emitting regions (the compact object and/or accretion disk). There is a species-dependence overlaid on this trend that has two principal origins. Firstly, charged-current processes involving nucleons generate imbalances that reflect the neutron-to-proton ratio. For example, in neutron-rich environments, neutrino absorption on nucleons delays the decoupling of $\nu_{e}$ relative to $\bar{\nu}_{e}$. Secondly, neutral-current scattering acts more strongly on $\bar{\nu}_{e}$ than $\nu_{e}$ due to the former’s higher average energy. These factors tend to deplete $\nu_{e}$ in outgoing directions and enhance $\bar{\nu}_{e}$ in ingoing directions. The result is that FFI is pervasive in multidimensional simulations, with an associated length scale commonly on the order of centimeters.
Evidence indicates that collisional flavor instability (CFI) [30] is also prevalent [31, 32, 33, 34, 35, 36]. This type of instability results from sufficiently distinct collision rates of $\nu_{e}$ and $\bar{\nu}_{e}$. The prime locations for CFI are the neutrino decoupling regions. In trapping regions, flavor instabilities of all kinds tend to be suppressed by the electron-flavor chemical potential $\mu_{\nu_{e}}$. In free-streaming regions, collisional rates are so low that CFI is probably unimportant even where it does appear. But in between, CFI is common and typically grows on a length scale of $\mathcal{O}(10)$ km. This scale can be much smaller when CFI is resonant [32, 37].
@@ 2.3 Estimated effects of oscillations
Figure 3: Potential astrophysical effects of neutrino oscillations. Left: Angle-averaged shock radius in 2D CCSN simulations with flavor conversion setting in above different critical densities (color-coded, with labels in g cm-3), showing significant deviations from the simulation without oscillations (“noFC”). Adapted with permission from Ehring et al. (2023) [51]. Middle: Nucleosynthesis abundance distributions formed in the simulated ejecta of a post-merger accretion disk, for various flavor-mixing prescriptions. Enhanced lanthanide abundances are observed. Modified with permission from Just et al. (2022) [52]. Right: CCSN neutrino energy spectra in various possible flavor-mixing scenarios, with potentially observable consequences. Adapted with permission from Abbar and Volpe (2024) [53].
@@
Flavor instabilities open the door to much more substantial effects on the astrophysics. For instance, as we saw in Sec. 2.2, flavor conversion likely occurs behind the supernova shock. Estimates indicate that explosion dynamics could be significantly affected [68, 51, 69, 70, 71]. The left panel of Fig. 3 exemplifies the findings obtained using approximate treatments of oscillations in CCSN simulations. Notably, the evolution of the shock is highly variable depending on precisely where the flavor conversion sets in. This observation suggests that rough approximations, though clearly valuable, are not an adequate answer to the oscillation problem. The details matter.
Flavor instabilities might augur significant changes to nucleosynthesis [72, 73, 28, 52, 74, 75, 76] (middle panel of Fig. 3). For example, flavor conversion in a CCSN may substantially affect neutrino-induced nucleosynthesis in outer stellar envelopes [64, 66]. A strong influence is also possible in the outflows from post-merger accretion disks, which are sensitive to neutrinos emanating from the accretion disk and the central neutron star if one exists. In this case, flavor conversion alters the electron fraction in the disk outflow and affects the yield of heavy neutron-rich elements [77, 28, 52, 29, 75, 76]. The modification of nuclide abundances could in turn appreciably affect the kilonova light curve produced by the decay of radioactive elements [28, 52]. Firming up the quantitative connection between flavor conversion and kilonovae is a high priority for this research area.
@@
The influence of neutrino flavor conversion on gravitational-wave emission is only just beginning to be assessed, but here again there are tantalizing hints that the impacts may be appreciable [79, 75]. In a CCSN, greater emission could result if, for example, protoneutron star convection is enhanced by flavor conversion in the neutrino decoupling region. In a NSM, it could result, for example, from an accelerated contraction of the outer layers of the remnant.
@@ 3.1 Quantum coherence and Wigner functions
However, solving for the complete quantum evolution is inconceivable from a practical standpoint. The spirit of quantum kinetics is to be selective in retaining and evolving information characterizing the field. Expectation values of the form $\langle a^{\dagger}(t)a(t)\rangle$ are generalized occupation numbers, with the generalization referring to the fact that $a^{\dagger}$ and $a$ may differ in their mass states, helicities, or momenta. In such cases, and in the chosen basis, $\langle a^{\dagger}(t)a(t)\rangle$ is not a density of particles per se but rather a density of quantum coherence. This review is concerned with mass coherence, or nonzero expectation values of the type $\langle a^{\dagger}_{j}(t)a_{i}(t)\rangle$ for $j\neq i$. Helicity coherence, $\langle a^{\dagger}_{h^{\prime}}(t)a_{h}(t)\rangle\neq 0$ for $h^{\prime}\neq h$, and opposite-momentum pairing, $\langle b(t,-\bm{q})a(t,\bm{q})\rangle\neq 0$, are generally suppressed by the high scale of neutrino kinetic energy [80, 81, 82, 83, 84].
Some amount of momentum coherence must be tolerated to allow for spatial inhomogeneity. To this end, the Wigner function (conventionally called the density matrix) is introduced [85, 86, 87]:
@@ 3.2 Coherent and incoherent processes
The commutator $H_{\bm{q}},\rho_{\bm{q}}]$ accounts for coherent processes [12, 13, 14, 15, 2]:
@@ 3.3 Many-body correlations
Several challenges will have to be met as the quantum many-body formulation of neutrino transport continues to develop. One is to find ways of tolerating a very large number of neutrinos given the exponential scaling of the dimension of Hilbert space [98, 99, 100, 101]. Another originates from the fact that Eq. 6 is actually a requantization of the kinetic Hamiltonian (Eq. 4) and, as such, unphysically privileges forward scattering [102, 103]. At the time of writing, only one study has carried out calculations using the complete many-body Hamiltonian [104]. It reported outcomes qualitatively different from those calculated using Eq. 6, observing flavor equilibration and momentum thermalization on the same time scale.
A final challenge concerns the issue of what initial states are appropriate for calculations. In particular, the applicability of quantum kinetics may hinge on the existence of localized particles [102, 103]. All fully many-body calculations performed so far have evolved states that are initially product states in the basis of definite-$\bm{q}$ plane waves. It is unclear whether kinetic evolution should be expected for such systems. The quasi-many-body model of Ref. [105] highlights this point. Two-body correlations build up during each brief interaction but are then traced out, following the logic of molecular chaos. The evolution in this model exhibits distinct coherent and incoherent contributions, as expected of a kinetic system. An important goal for this research area is to definitively establish the relationship between kinetic evolution and the properties of the underlying quantum state or ensemble.
@@ 4.1 Symmetries and conservation laws
An advantage of polarization vectors is that they draw out the connection between neutrino flavor and angular momentum. Consider the QKE for $\bm{P}_{\bm{q}}$, which follows from plugging the definition above into Eq. 3:
@@
with $\omega_{\bm{q}}=\delta m^{2}/2|\bm{q}|$ and $\delta m^{2}=m_{2}^{2}-m_{1}^{2}$. The signs of $\delta m^{2}$ and $\omega_{\bm{q}}$ reflect the neutrino mass ordering. All flavor-space vectors are defined through expansions similar to Eq. 8. We write $\bm{\vec{D}}_{1}$ with an arrow to distinguish it from $\bm{D}_{1}$ in Eq. 5. By our chosen $\bar{\rho}$ convention, $\bm{\bar{H}}_{\bm{q}}$ is identical to $\bm{H}_{\bm{q}}$ but with $\omega\bm{B}\rightarrow-\omega\bm{B}$. We will stick with the flavor basis from this point on, so that $\bm{z}$ always represents the flavor axis. Assuming the muon and tau populations are negligible, the charged-lepton vector is $\bm{L}=n_{e}\bm{z}$, with net electron rest number density $n_{e}$. The mass vector is $\bm{B}=(\sin 2\theta)\bm{x}-(\cos 2\theta)\bm{z}$, where $\theta$ is the vacuum mixing angle.
@@
A defining feature of the oscillation problem is that calculations are only tractable for relatively simple models. Simplification, however, is an art. Often it introduces symmetries and additional invariants that can artificially stabilize neutrino flavor states and otherwise fundamentally change the dynamics, as strikingly demonstrated by the slow [110, 111, 112, 113], fast [114, 115, 112, 113], and collisional [116] flavor pendula. To borrow an adage from Nigel Goldenfeld: Symmetry is the enemy of instability. The potential for insight is balanced by the risk of unwarranted extrapolation. With all models, it is wise to ask which features are general, which are informatively specific, and which are just irrelevant.
@@ 4.2 Adiabaticity
One of the puzzles to come out of the bulb model was that adiabaticity sometimes accounted for the numerically observed behavior and sometimes conflicted with it. The effort to determine the conditions under which oscillations are adiabatic motivated the formalization of linear stability analysis as it applies to neutrino flavor (Sec. 4.3). Ultimately, the systematic study of flavor instabilities did not merely clarify the evolution that had been observed numerically. It undermined the use of the bulb model as a way to predict flavor transformation in the first place. The reason was already alluded to in Sec. 4.1: the stringent assumptions of the bulb model restrict the evolution in unacceptable ways, as evidenced by the presence of symmetry-breaking flavor instabilities. The story for NSMs parallels the one for CCSNe. In simple merger models, adiabaticity explains the numerically observed evolution through matter–neutrino resonances [59, 60, 123, 124, 125, 126]. These models appear to be compromised by oversimplification as well [127, 128].
@@ 4.3 Instability
By itself, this procedure does not shed much light on what gives rise to flavor instabilities. Deeper understanding comes from considering the linearized evolution of a neutrino system in certain idealized limits. Here the common three-type classification of instabilities enters (Sec. 2.2). The first class is FFI [20, 21]. By definition, an instability is fast if it occurs when all vacuum oscillation frequencies $\omega$ and all collision rates $\Gamma$ are set to zero in the dispersion relation. Growth rates can be as high as $\mathcal{O}(\mu)$, where $\mu=\sqrt{2}G_{F}n_{\nu}$. The category of SFI includes all the collisionless instabilities that disappear at $\omega=0$ [38, 39]. The speediest among them have $\mathcal{O}(\sqrt{\mu\omega})$ growth rates. CFI rounds out the taxonomy [30]. As the name implies, CFI disappears at $\Gamma=0$. Its growth rates reach up to $\mathcal{O}(\sqrt{\mu\Gamma})$. It is important to stress that these rates are highly approximate and are not even the correct scaling in all cases. For example, FFI identified in a simulation usually has a rate quite a bit smaller than $\mu$, and CFI more typically has a rate roughly proportional to $\Gamma$.
Occasionally one encounters other instability types, such as bipolar, multi-azimuthal-angle, and multi-zenith-angle. These terms are somewhat more historical, though they can still be useful depending on the context. For the newcomer to collective oscillations, the important thing to know is that they do not refer to distinct phenomena not covered here.
@@
The concept of self-induced resonance helps illuminate the physics of flavor instabilities [138, 114]. We will explain the basic idea using the equation of motion $\partial_{t}\bm{P}_{\bm{q}}=\bm{H}_{\bm{q}}\times\bm{P}_{\bm{q}}$. Each polarization vector $\bm{P}_{\bm{q}}$ attempts to precess around its Hamiltonian $\bm{H}_{\bm{q}}$, but $\bm{H}_{\bm{q}}$ itself is evolving. If the relative azimuthal angle between these vectors evolves rapidly, then the term $\bm{H}_{\bm{q},T}\times\bm{P}_{\bm{q},T}$, which drives the evolution of $P_{\bm{q},z}$, is expected to average to zero over time. However, if $\bm{H}_{\bm{q},T}$ and $\bm{P}_{\bm{q},T}$ maintain a fixed relative phase over a long enough duration—in other words, if there is a resonance between the frequencies—then a secular change in $P_{\bm{q},z}$ can accumulate. This idea extends to spatially inhomogeneous systems, where individual neutrinos can additionally be resonant with $\bm{K}\neq 0$ collective waves [132]. This is essentially the same particle–wave resonance mechanism that is central to Landau damping in plasmas and other physical systems.
Resonance functions somewhat differently in CFI [30, 116]. Consider the idealized system governed by $\partial_{t}\bm{P}=\sqrt{2}G_{F}\bm{D}\times\bm{P}-\Gamma\bm{P}_{T}$ and $\partial_{t}\bm{\bar{P}}=\sqrt{2}G_{F}\bm{D}\times\bm{\bar{P}}-\bar{\Gamma}\bm{\bar{P}}_{T}$, which are coupled through $\bm{D}=\bm{P}-\bm{\bar{P}}$. Introducing $\Gamma^{\pm}=(\Gamma\pm\bar{\Gamma})/2$, we rewrite these equations in terms of the difference vector $\bm{D}$ and sum vector $\bm{S}=\bm{P}+\bm{\bar{P}}$: